# ROTATE: Regret-driven Open-ended Training for Ad Hoc Teamwork

### v1

@@ Appendix C Supplemental Results
This section presents various supplemental results. First, we describe CoMeDi’s mixed-play mechanism in the context of ROTATE’s per-state regret. Second, we discuss alternative estimators for ROTATE per-state regret. Third, we present experiments comparing ROTATE to a variant with CoMeDi-style mixed-play return maximization, and a variant using the alternative regret estimation strategy. Next, we present and describe radar charts breaking down the performance of ROTATE on all six tasks presented in the main paper. Finally, we present the learning curves for all variants of ROTATE that are tested in this paper.
@@ C.1 Discussion of CoMeDi and Mixed Play
[…]
@@
Despite the importance of using $p_{\text{MSTART}}$ as a starting state for data collection being questionable, we take inspiration from CoMeDi’s maximization of $V(s|\pi,\text{BR}(\pi))$ outside of states from $p_{\text{SP}}$. We argue that maximizing $V(s|\pi^{-i},\text{BR}(\pi^{-i}))$ is a key component towards making $\pi^{-i}$ act in good faith by always choosing actions yielding optimal collective returns assuming $\text{BR}(\pi^{-i})$ is substituted as the partner policy. Unlike CoMeDi, ROTATE maximizes $V(s|\pi^{-i},\text{BR}(\pi^{-i}))$ on trajectories gathered from a starting state from $p_{\text{XP}}$ instead of $p_{\text{MSTART}}$, which results in the second term of Expr. 10. We formulate this objective to encourage $\pi^{-i}$ to act in good faith in states sampled from $p_{\text{XP}}$, which is visited while $\pi^{-i}$ interacts with $\pi^{\text{ego}}$. Since $\pi^{-i}$ is not sabotaging $\pi^{\text{ego}}$ by selecting actions that make collaboration impossible in $p_{\text{XP}}$, the ego policy learning process becomes less challenging. We conjecture that this leads to $\pi^{\text{ego}}$ with better performances as indicated in Figure 3.
@@ C.2 Alternatives Estimators for Per-State Regret
20 in Algorithm 2 and Eq. 11 outline how ROTATE maximizes per-state regret in states visited during XP interaction (denoted by $p_{\text{XP}}$), where SP and XP returns are estimated via a trained critic and a 1-step return estimate, respectively. As a reminder, ROTATE employs the following target function to train the regret-maximizing teammate policy:
@@
We maximize regret in states sampled from $p_{\text{XP}}$ to encourage the design of teammate policies that provide a learning challenge while also acting in good faith, thereby maximizing cooperative returns assuming interactions with its best-response policy, while interacting with the ego agent’s policy. Despite potentially providing biased estimates, training a value function to estimate self-play returns can reduce the variance caused by environment stochasticity, compared to a Monte Carlo return-to-go estimate.
The critic network estimating teammate-BR returns, $V_{\sigma^{-i,\text{BR}}}(s)$, is trained on interactions initialized from states sampled from $p_{\text{XP}}$ (SXP interactions), as shown in 22 of Algorithm 2. This enables the teammate-BR critic network to accurately estimate SP returns from $p_{\text{XP}}$ states. Meanwhile, a 1-step estimate of XP returns is made possible by storage of rewards experienced during XP interactions (12 of Algorithm 2) and the training of a value function to estimate XP returns (23 of Algorithm 2). Utilizing a 1-step estimate produces lower variance than using a Monte Carlo-based return-to-go estimate, while also yielding less bias than predicting returns solely based on the trained critic network’s value.
@@
$\mathbb{E}_{s_{t}\sim p_{\text{XP}}}\left[\underbrace{\mathbb{E}_{\begin{subarray}{c}a_{t^{\prime}}\sim[\text{BR}(\pi^{-i}),\pi^{-i}],P\end{subarray}}\Bigg{[}\sum_{t^{\prime}=t}^{\infty}\gamma^{t^{\prime}}r_{t^{\prime}}\Bigg{|}s_{t}\Bigg{]}}_{\text{SP return estimate}}-\>\>\underbrace{\mathbb{E}_{\begin{subarray}{c}a_{t^{\prime}}\sim[\pi^{\text{ego}},\pi^{-i}],P\end{subarray}}\Bigg{[}\sum_{l=0}^{\infty}\gamma^{t^{\prime}}r_{t^{\prime}}\Bigg{|}s_{t}\Bigg{]}}_{\text{XP return estimate}}\right].$ (20)
@@
$\mathbb{E}_{s_{t}\sim p_{\text{XP}}}\left[\underbrace{\mathbb{E}_{\begin{subarray}{c}a_{t^{\prime}}\sim[\text{BR}(\pi^{-i}),\pi^{-i}],P\end{subarray}}\Bigg{[}\underbrace{\sum_{t^{\prime}=t}^{\infty}(\gamma\lambda)^{t^{\prime}}\>\delta^{-i,\text{BR}}_{t^{\prime}}}_{\text{GAE}}\Bigg{|}s_{0}\Bigg{]}}_{\text{SP return estimate}}-\>\>\underbrace{\mathbb{E}_{\begin{subarray}{c}a_{t^{\prime}}\sim[\pi^{\text{ego}},\pi^{-i}],P\end{subarray}}\Bigg{[}\underbrace{\sum_{t^{\prime}=t}^{\infty}(\gamma\lambda)^{t^{\prime}}\>\delta^{-i,\text{ego}}_{t^{\prime}}}_{\text{GAE}}\Bigg{|}s_{0}\Bigg{]}}_{\text{XP return estimate}}\right],$ (21)
@@ C.3 Experimental Comparisons of ROTATE Teammate Generation Objectives
(b) ROTATE ego agent vs independently trained ego agent on ROTATE’s population.
@@
C.4 ROTATE vs Baselines—Radar Charts
@@ Appendix D Experimental Tasks
Experiments in the main paper are conducted on Jax re-implementations of Level-Based Foraging (LBF) [2, 5], and five tasks from the Overcooked suite—Cramped Room (CR), Asymmetric Advantages (AA), Counter Circuit (CC), Coordination Ring (CoR), and Forced Coordination (FC) [8, 45]. Each task is described below.
@@
The objective for all five tasks is to deliver as many onion soups as possible, where the only difference between the tasks is the environment layout, as shown in Fig. 9. To deliver an onion soup, players must place three onions in a pot to cook, use a plate to pick up the cooked soup, and send the plated soup to the delivery location. Each player observes the state and location of all environment features (counters, pots, delivery, onions, and plates), the position and orientation of both players, and an urgency indicator, which is 1 if there are 40 or fewer remaining time steps, and 0 otherwise. Each player has six discrete actions, consisting of the four movement actions, interact, and no-op. The reward function awards both agents $+20$ upon successfully delivering a dish, which is the return reported in the experimental results. To improve sample efficiency, all algorithms are trained using a shaped reward function that provides each agent an additional reward of $0.1$ for picking up an onion, $0.5$ for placing an onion in the pot, $0.1$ for picking up a plate, and $1.0$ for picking up a soup from the pot with a plate. An episode terminates after 400 time steps. Player locations are randomized in each episode. In divided layouts such as AA and FC, we ensure that a player is spawned on each half of the layout.
@@ Appendix E Implementation Details
As implementations of prior methods use PyTorch, but this project uses Jax, we re-implemented all methods in this paper, using PPO [48] with Generalized Advantage Estimation (GAE) [47] as a base RL algorithm and Adam [26] as the default optimizer. The code for this paper is released at https://github.com/carolinewang01/rotate, and we recommend consulting it for a full understanding of method implementations. Pseudocode for ROTATE is provided in App. A. This section discusses implementation details such as training time choices, agent architectures, and key hyperparameters for ROTATE and all baselines.
@@ E.2 Agent Architectures
In particular, for ROTATE and PAIRED, the teammate must estimate returns when interacting with the ego agent and its best response, and so it maintains a critic network for each partner type. For CoMeDi and BRDiv, given a population with $n$ agents, each teammate must estimate the return when interacting with the other $n-1$ agents in the population. As it would be impractical to maintain $n-1$ critics for each teammate, the teammate instead uses a critic that conditions on the agent ID of a candidate partner agent—in effect, implementing the $n-1$ critics via parameter sharing [12].
Task LBF CR AA CC CoR FC
@@ E.3 Hyperparameters
This section presents the hyperparameters for ROTATE (Table 3), baseline methods (Tables 2, 4, 5, 6, 7 and 8), and training evaluation teammates with IPPO (Table 1). Note that hyperparameters for the two-stage teammate generation methods are presented in separate tables, where those corresponding to the shared ego agent training stage are presented in Table 4. All experiments in the paper were performed with a discount factor of $\gamma=0.99$ and $\lambda^{\text{GAE}}=0.95$.
@@ Appendix F Evaluation Teammate Details
Descriptions of the evaluation teammates for each task and estimated best response returns are provided in Tables 9, 10, 11, 12, 13 and 14.

### v2

@@ Appendix C Supplemental Results
This section presents various supplemental results. First, we describe CoMeDi’s mixed-play mechanism in the context of ROTATE’s per-state regret. Second, we discuss alternative estimators for ROTATE per-state regret. Third, we present experiments comparing ROTATE to a variant with CoMeDi-style mixed-play return maximization, and a variant using the alternative regret estimation strategy. Fourth, we examine whether the population generated by ROTATE is useful for training an independent ego agent. Fifth, we present and describe radar charts breaking down the performance of ROTATE on all six benchmark tasks presented in the main paper, and the learning curves for all variants of ROTATE that are tested in this paper. Finally, we present a human proxy evaluation of ROTATE and CoMeDi on the Overcooked tasks.
@@ C.1 Discussion of CoMeDi and Mixed Play
$\underset{\pi}{\max\ }(\mathbb{E}_{s_{0}\sim p_{0}}\left[\text{CR}(\pi^{\text{comp}},\pi^{-i},s_{0})\right]+\underbrace{\mathbb{E}_{s\sim p_{\text{MP}}}[V(s|\pi,\text{BR}(\pi))]}_{\text{mixed-play return maximization}}).$ (16)
@@
Despite the importance of using $p_{\text{MSTART}}$ as a starting state for data collection being questionable, we take inspiration from CoMeDi’s maximization of $V(s|\pi,\text{BR}(\pi))$ outside of states from $p_{\text{SP}}$. We argue that maximizing $V(s|\pi^{-i},\text{BR}(\pi^{-i}))$ is a key component towards making $\pi^{-i}$ act in good faith by always choosing actions yielding optimal collective returns assuming $\text{BR}(\pi^{-i})$ is substituted as the partner policy. Unlike CoMeDi, ROTATE maximizes $V(s|\pi^{-i},\text{BR}(\pi^{-i}))$ on trajectories gathered from a starting state from $p_{\text{XP}}$ (i.e., SXP states) instead of $p_{\text{MSTART}}$, which results in the second term of Expr. 10. We formulate this objective to encourage $\pi^{-i}$ to act in good faith in states sampled from $p_{\text{XP}}$, which is visited while $\pi^{-i}$ interacts with $\pi^{\text{ego}}$. Since $\pi^{-i}$ is not sabotaging $\pi^{\text{ego}}$ by selecting actions that make collaboration impossible in $p_{\text{XP}}$, the ego policy learning process becomes less challenging. We conjecture that this leads to $\pi^{\text{ego}}$ with better performances as indicated in Figure 3.
@@ C.2 Alternatives Estimators for Per-State Regret
˜21 in Algorithm 2 and Eq. 11 outline how ROTATE maximizes per-state regret in states visited during XP interaction (denoted by $p_{\text{XP}}$), where SP and XP returns are estimated via a trained critic and a 1-step return estimate, respectively. As a reminder, ROTATE employs the following regret target function to train the regret-maximizing teammate policy on XP states, with an analogously defined target function for SP states:
@@
We maximize regret in states sampled from $p_{\text{XP}}$ and $p_{\text{SP}}$ to encourage the design of teammate policies that provide a learning challenge while also acting in good faith, thereby maximizing cooperative returns assuming interactions with its best-response policy, while interacting with the ego agent’s policy. Our discussion here focuses on computing Eq. 19 for brevity. However, a similar approach can be used to train a critic network to estimate regret in SP states accurately. The only difference lies in the use of states sampled from $p_{\text{SP}}$ and $p_{\text{XSP}}$ for training the critic network.
Despite potentially providing biased estimates, training a value function to estimate self-play returns can reduce the variance caused by environment stochasticity, compared to a Monte Carlo return-to-go estimate.
The critic network estimating teammate-BR returns, $V_{\sigma^{-i,\text{BR}}}(s)$, is trained on interactions initialized from XP, as shown in ˜23 of Algorithm 2. This enables the teammate-BR critic network to accurately estimate SP returns from $p_{\text{XP}}$ states. Meanwhile, a 1-step estimate of XP returns is made possible by storage of rewards experienced during XP interactions (˜12 of Algorithm 2) and the training of a value function to estimate XP returns (˜24 of Algorithm 2). Utilizing a 1-step estimate produces lower variance than using a Monte Carlo-based return-to-go estimate, while also yielding less bias than predicting returns solely based on the trained critic network’s value.
@@
$\mathbb{E}_{s_{t}\sim 0.5p_{\text{XP}}+0.5p_{\text{SP}}}\left[\underbrace{\mathbb{E}_{\begin{subarray}{c}a_{t^{\prime}}\sim[\text{BR}(\pi^{-i}),\pi^{-i}],P\end{subarray}}\Bigg[\sum_{t^{\prime}=t}^{\infty}\gamma^{t^{\prime}}r_{t^{\prime}}\Bigg|s_{t}\Bigg]}_{\text{SP return estimate}}-\>\>\underbrace{\mathbb{E}_{\begin{subarray}{c}a_{t^{\prime}}\sim[\pi^{\text{ego}},\pi^{-i}],P\end{subarray}}\Bigg[\sum_{l=0}^{\infty}\gamma^{t^{\prime}}r_{t^{\prime}}\Bigg|s_{t}\Bigg]}_{\text{XP return estimate}}\right].$ (20)
@@
$\mathbb{E}_{s_{t}\sim 0.5p_{\text{XP}}+0.5p_{\text{SP}}}\left[\underbrace{\mathbb{E}_{\begin{subarray}{c}a_{t^{\prime}}\sim[\text{BR}(\pi^{-i}),\pi^{-i}],P\end{subarray}}\Bigg[\underbrace{\sum_{t^{\prime}=t}^{\infty}(\gamma\lambda)^{t^{\prime}}\>\delta^{-i,\text{BR}}_{t^{\prime}}}_{\text{GAE}}\Bigg|s_{0}\Bigg]}_{\text{SP return estimate}}-\>\>\underbrace{\mathbb{E}_{\begin{subarray}{c}a_{t^{\prime}}\sim[\pi^{\text{ego}},\pi^{-i}],P\end{subarray}}\Bigg[\underbrace{\sum_{t^{\prime}=t}^{\infty}(\gamma\lambda)^{t^{\prime}}\>\delta^{-i,\text{ego}}_{t^{\prime}}}_{\text{GAE}}\Bigg|s_{0}\Bigg]}_{\text{XP return estimate}}\right],$ (21)
@@ C.3 Experimental Comparisons of ROTATE Teammate Generation Objectives
(b) ROTATE compared to an independently trained ego agent on ROTATE’s population and an ablation of ROTATE without the population.
@@
C.4 Training an Independent Ego Agent on the ROTATE Population
Two-stage AHT algorithms first generate a population of teammates, and next train an ego agent against the population. Although ROTATE’s teammate generation mechanism relies on the learning process of a particular ego agent, here, we investigate whether the population generated by ROTATE is useful for training independently generated ego agents. Fig. 6(b) compares the mean evaluation returns of the ROTATE ego agent against the mean evaluation returns of an independently trained ego agent that was trained using the same configuration as ROTATE. In 3/6 tasks, the ROTATE ego agent outperforms the trained ego agent, while in two tasks, the two ego agents perform similarly (LBF and FC). This result suggests that the ROTATE population is a useful population of teammates even independent of the particular ego agent generated. The strong performance of the independently trained ego agent is unsurprising given that it has two advantages over the ROTATE ego agent. First, the independently trained ego agent faces a stationary distribution of training teammates compared to ROTATE, which faces a nonstationary distribution caused by the population growing over learning iterations. Second, the independently trained ego agent interacts with all teammates uniformly throughout training, while the ROTATE ego agent only trains against earlier teammates for more iterations than later teammates.
C.5 ROTATE vs Baselines—Radar Charts
@@ C.6 Learning Curves
C.7 Human Proxy Evaluations
The results in the main paper show that ROTATE is better able to generalize to unseen partners compared to baseline methods. Here, we evaluate the ability of ROTATE to generalize to human partners, compared to the best baseline, CoMeDi. The evaluation is conducted with human proxy teammates generated by behavior cloning on the human gameplay dataset published by [9]. Table 2 shows that ROTATE achieves higher returns in coordination with the human proxy compared to CoMeDi, across all five Overcooked layouts.
CR AA $C$ CoR FC
ROTATE 0.85 (0.82, 0.87) 0.68 (0.65, 0.70) 0.92 (0.88, 0.95) 0.66 (0.65, 0.68) 0.60 (0.56, 0.64)
CoMeDi 0.75 (0.71, 0.78) 0.57 (0.54, 0.59) 0.89 (0.83, 0.95) 0.56 (0.54, 0.59) 0.40 (0.38, 0.43)
Table 2: ROTATE outperforms CoMeDi with the human proxy teammate on all Overcooked layouts. Normalized returns and bootstrapped 95% CI’s are shown, where the normalization is performed using the human proxy agent’s self-play returns.
@@ Appendix D Experimental Tasks
Experiments in the main paper are conducted on Jax re-implementations of Level-Based Foraging (LBF) [2, 6], five tasks from the Overcooked suite (Cramped Room (CR), Asymmetric Advantages (AA), Counter Circuit (CC), Coordination Ring (CoR), and Forced Coordination (FC)) [9, 45] and a sabotage matrix game. The sabotage matrix game is already fully described in the main paper. The other tasks are described below.
@@
The objective for all five tasks is to deliver as many onion soups as possible, where the only difference between the tasks is the environment layout, as shown in Fig. 10. To deliver an onion soup, players must place three onions in a pot to cook, use a plate to pick up the cooked soup, and send the plated soup to the delivery location. Each player observes the state and location of all environment features (counters, pots, delivery, onions, and plates), the position and orientation of both players, and an urgency indicator, which is 1 if there are 40 or fewer remaining time steps, and 0 otherwise. Each player has six discrete actions, consisting of the four movement actions, interact, and no-op. The reward function awards both agents +20+20 upon successfully delivering a dish, which is the return reported in the experimental results. To improve sample efficiency, all algorithms are trained using a shaped reward function that provides each agent an additional reward of $0.1$ for picking up an onion, $0.5$ for placing an onion in the pot, $0.1$ for picking up a plate, and $1.0$ for picking up a soup from the pot with a plate. An episode terminates after 400 time steps. Player locations are randomized in each episode. In divided layouts such as AA and FC, we ensure that a player is spawned on each half of the layout.
@@ Appendix E Implementation Details
As implementations of prior methods use PyTorch, but this project uses Jax, we re-implemented all methods in this paper, using PPO [48] with Generalized Advantage Estimation (GAE) [47] as a base RL algorithm and Adam [26] as the default optimizer. The code for this paper is released for reproducibility at https://github.com/carolinewang01/rotate. Pseudocode for ROTATE is provided in App. A. This section discusses implementation details such as training time choices, agent architectures, and key hyperparameters for ROTATE and all baselines.
@@ E.2 Agent Architectures
In particular, for ROTATE and PAIRED, the teammate must estimate returns when interacting with the ego agent and its best response, and so it maintains a critic network for each partner type. For CoMeDi and BRDiv, given a population with $n$ agents, each teammate must estimate the return when interacting with the other n1n-1 agents in the population. As it would be impractical to maintain n1n-1 critics for each teammate, the teammate instead uses a critic that conditions on the agent ID of a candidate partner agent—in effect, implementing the n1n-1 critics via parameter sharing [13].
Task LBF CR AA $C$ CoR FC
@@ E.3 Hyperparameters
This section presents the hyperparameters for ROTATE (Table˜5), baseline methods (Tables˜4, 6, 7, 8, 9 and 10), and training evaluation teammates with IPPO (Table˜3). Note that hyperparameters for the two-stage teammate generation methods are presented in separate tables, where those corresponding to the shared ego agent training stage are presented in Table˜6. All experiments in the paper were performed with a discount factor of $\gamma=0.99$ and $\lambda^{\text{GAE}}=0.95$.
@@ Appendix F Evaluation Teammate Details
Descriptions of the evaluation teammates for each task and estimated best response returns are provided in Tables˜11, 12, 13, 14, 15 and 16.